OpenStudy (anonymous):

(-1, 3) and perpendicular to y = -negative one thirdx + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

the line \[y=-\frac{1}{3}x+7\] has slope \[-\frac{1}{3}\] so a perpendicular line will have slope 3

OpenStudy (anonymous):

\[y-3=3(x+1)\] \[y-3=3x+3\] \[y=3x+6\]